If the fitness of AA, Aa, and aa are 1.0, 0.9, 0.6, and p0 = 0.7, calculate p1, p2, and p3, the allele frequencies after 1, 2, and 3 generations of selection.
Answer #1

Given

Po = 0.7

Qo = 1-Po = 1-0.7 = 0.3

WAA = 1.0

WAa = 0.9

Waa = 0.6

P1 = Po(PoWAA + QoWAa) / W'

W' = Average fitness = Po2 WAA + 2PoQoWAa + Qo2 Waa

W' = (0.7*0.7*1) + (2*0.7*0.3*0.9) + (0.3*0.3*0.6) = 0.922

P1 = 0.7(0.7+(0.3*0.9))/0.922 = 0.74

Q1 = 0.26

Similarly calculate for P2 and P3.

P2 = 0.74 ( 0.74 + (0.26*0.9))/0.922 = 0.78

Q2 = 0.22

P3 = 0.78(0.78 + (0.22*0.9) )/ 0.922 = 0.827

Q3 = 0.173